OpenStudy (anonymous):

Find the sum of the infinite geometric series: 9 – 6 + 4 – …

OpenStudy (raden):

use formula : S = a/(1-r) for a = the first term r = ratio

OpenStudy (anonymous):

The thing is I don't know how to find r

OpenStudy (raden):

r = a2/a1 = -6/9 = -2/3

OpenStudy (raden):

so, S = 9/(1-(-2/3)) = 9/(1+2/3) = 9/(5/3) = ... ?

OpenStudy (raden):

or u can write into multiplication : 9 * (3/5) = ...